# Effective programming of a photonic processor with complex interferometric structure

### v1

@@ Appendix E Unitaries on chip measurement
where 2N1=72N-1=7 values of $\tilde{\gamma}_{ij}$ have zero values by construction. These are the angles of the $l_{1}$-th column and $l_{1}$-th row of the reconstructed matrix $\tilde{U}$ on the chip. The fact that in the reconstructed unitary transformation $\tilde{U}$ of the target $N$-port interferometer, 2N12N-1 angles are set to be real is inherent to any tomography algorithm [47, 43, 44, 48]. This is because of the physical equivalence of $U$ and $\tilde{U}=P_{out}(\vec{\varphi_{out}})\times U\times P_{in}(\vec{\varphi_{in}})$, where $P_{in}(\vec{\varphi_{in}})$ and $P_{out}(\vec{\varphi_{out}})$ are diagonal matrices with pure phases $\vec{\varphi_{in}}$ and $\vec{\varphi_{out}}$, respectively. In other words, any tomography algorithm can reconstruct the unitary transformation of a linear optical interferometer only up to the input and output phases [47, 43, 44, 48].
As a final remark on the reconstruction procedure, the modules and angles measurements are conducted completely independently, which leaves no guarantee that the reconstructed by (23) matrix $\tilde{U}$ is unitary. Even in case if one had performed the Sinkhorn Knopp algorithm [46] to rescale the experimentally measured modules of $\tilde{U}$, the angles were still obtained using a separate interferometric procedure. However, if $\tilde{U}$ appeared to be a non unitary, it can be projected to the closest matrix from the unitary subspace by, for example, polar decomposition or single value decomposition [49].
As, typically, optical power from the all output ports of the chip is being collected simultaneously each port by each photodetector, the number of separate measurements necessary for obtaining the modules $|U_{ij}|$ for an $N\times N$-port interferometr is equal to $N$. Whereas the angles $\tilde{\gamma}_{ij}$ measurement requires N1N-1 separate interference collections, resulting in 2N12N-1 total number of measurements for the target $N\times N$ unitary reconstruction.
In our experiment, we set $l_{1}=1$ for every measurement of $\tilde{U}$ on the chip, which causes the first row and first column of any measured $\tilde{U}$ to be strictly real. However, because the measured unitary matrix of the target interferometer $\tilde{U}$ is always in the form of a real first row and first column, the simulated $U_{sim}$ matrix, generated by the chips’ digital model, should be converted to this same form before their comparison, as the matrix fidelity (9) measure is sensitive to such deviations, while these unitaries are physically equivalent. This type of unitary matrix $U_{sim}=\rho_{ij}\exp(i\varphi_{ij})$ transformation can easily be performed by multiplying it on both sides by diagonal unitary matrices:
@@
where $\Gamma=diag(e^{i\gamma_{j}})$ and $\Omega=diag(e^{i\omega_{i}})$, the elements of which are equal to $\gamma_{j}=-\varphi_{1j}+\varphi_{11}$ and $\omega_{i}=-\varphi_{i1}$. After multiplying by $\Gamma$ and $\Omega$ unitary matrix $U_{sim}$ transforms to $\tilde{U}_{sim}=\rho_{ij}\exp(i\tilde{\varphi}_{ij})$ with zero angles in the first row and first column by construction: $\tilde{\varphi}_{1j}\equiv\tilde{\varphi}_{i1}\equiv 0$.
More generally, any $N\times N$ unitary matrix $U_{ij}=\rho_{ij}\exp(i\varphi_{ij})$ can always be multiplied on both sides by a diagonal $N\times N$ unitary matrices $\Gamma=\text{diag}(e^{i\gamma_{j}})$ and $\Omega=\text{diag}(e^{i\omega_{i}})$ with appropriate angles $\gamma_{j}$ and $\omega_{i}$ such that the resulting unitary matrix $\tilde{U}_{ij}=\rho_{ij}\exp(i\tilde{\varphi}_{ij})$ has purely real $I$-th row and $J$-th column: $\tilde{\varphi}_{iJ}\equiv\tilde{\varphi}_{Ij}\equiv 0$. The necessary angles, $\gamma_{j}$ and $\omega_{i}$ are obtained from:
@@
Appendix F Broadband optical switching details
@@
Optical switches were realized via the chips’ model at several wavelengths ranging from $915$ to $975$ nm. At each wavelength, the digital chip model was reconstructed using the calibration method. The reconstructed mode mixing matrices $M_{1,2}$, bias phase shift vectors $\vec{\Phi}_{0}$ and crosstalk matrices $A$ for each wavelength are shown in Fig. 11.
@@
As shown in Fig. 11 the mode mixing matrices $M^{\lambda}_{1,2}$ are slightly different for each wavelength $\lambda$ preserving the property $M^{\lambda}_{1}\approx M^{\lambda}_{2}$. The bias phase shift vectors $\vec{\Phi}_{0}$ varies with wavelength $\lambda$ more significantly, but always in accordance with the actual geometry of the target interferometer: $\varphi_{1}^{\lambda}\approx 0$ and $\varphi_{2}^{\lambda}\approx\varphi_{3}^{\lambda}$ as discussed in Appendix D. Meanwhile, the crosstalk matrices $A^{\lambda}$ for each wavelength $\lambda$ are approximately the same. This is because the crosstalk matrix $A$ describes the physical heaters (and is completely determined by them) and not the coherent radiation.
@@
The necessary three phase shifts for implementing a particular port-to-port optical switch were obtained by an optimization procedure run on a standard PC (Python scipy.minimize method). Once the required phase shifts were determined from the digital model of the optical chip, they were applied to the actual device. Coherent radiation of a given wavelength was injected into the specific input port of the device. The optical power distribution at the output ports was then measured using photodiodes and compared with the target values. A comparison between the switching matrices $U$ and $V$ was performed according to the variation in the amplitude fidelity measure:
@@
An example simulating coherent light with a wavelength of $975$ nanometers propagating through a target interferometer, illustrating all possible port-to-port optical switch configurations, is shown in Fig. 12.
Appendix G Scaling of the proposed calibration procedure
The scalability of the proposed approach for programming a reconfigurable integrated interferometer is a crucial consideration that needs to be carefully addressed. First, in our experiment, the tunable $4\times 4$ integrated interferometer consists of two passive mode mixing blocks $M_{1,2}$ and an active phase shift layer $P_{1}(\vec{\varphi}_{1})$ between them. The phase shift layer consists of three thermo phase shifters. Therefore, our optical chip has only three tunable parameters and cannot be considered a $4\times 4$ universal photonic device. However, if two additional phase shift layers $P_{2}(\vec{\varphi}_{2})$ and $P_{3}(\vec{\varphi}_{3})$ were cascaded between the two additional constant mode mixing blocks $M_{3}$ and $M_{4}$ and added to our device, the resulting interferometer would become a $4\times 4$ universal interferometer [9, 31, 10].
@@ Appendix G Scaling of the proposed calibration procedure
This calibration will result in the exact phase from the current dependence for the first phase shift layer, as well as in the explicit form of the first mode mixing block $M_{1}$ and effective mode mixing block $M_{2}$:
@@
This calibration will result in the exact phase from the current dependence for the second phase shift layer, as well as in the explicit form of the third mode mixing block $M_{3}$ and effective mode mixing block $M_{1}$:
@@
which, again, just as $\tilde{M}_{2}$ implicitly contains the information about the actual $M_{2}$. However, we know the explicit forms of $M_{1}$ and $P_{1}(\vec{\varphi}_{1}^{(0)})$ (which is the first phase shift layer with zero currents). This enables us to determine the actual second mode mixing block $M_{2}$ as follows:
@@
The expressions (31) and (32) for calculating the $M_{2}$ are independent, as entities $M_{1}$, $P_{1}^{(0)}$, $\tilde{M}_{2}$ and $M_{3}$, $P_{2}^{(0)}$, $\tilde{M}_{1}$ were obtained from two independent experimental calibration procedures for the first (29) and second (30) phase shift layers $P_{1}(\vec{\varphi}_{1}(\vec{X}_{1}))$ and $P_{2}(\vec{\varphi}_{2}(\vec{X}_{2}))$, respectively. The two independent evaluations of $M_{2}$ must be equivalent and can be used to validate the entire optical chip digital model. Figure 9 schematically illustrates the procedure described above for programming an interferometer with two phase shift layers.
Figure 9: A model of an interferometer with two phase shift layers and three mode mixing layers can be consequently treated as a) interferometer with one phase shift layer cascaded between two mode mixing layers and b) interferometer with one phase shift layer cascaded between two mode mixing layers. The mode mixing block, $M_{2}$ can be implicitly determined.
@@
The required number of measurements for calibrating phase shift layer in a general $N\times N$ interferometer is $N$ input ports $\times N-1$ heaters, which is quadratic with respect to the number of modes. The number of real parameters required for optimization is also quadratic in the number of modes.
In other words, there are two ways to scale an interferometer: by increasing its length, or by increasing the width or the number of modes it contains. The first method is simpler and involves a linear increase in the number of measurements and amount of processing required. However, scaling by increasing the number of modes grows quadratically with both the number of measurements and parameters of the model that need to be optimized.
While the number of measurements cannot be decreased by increasing the number of modes, $N$, the number of parameters in the model can be reduced by carefully adjusting the model. Therefore, finding possible methods to reduce the number of model parameters is of great importance. For example, using prior knowledge of the structure of actual interferometers and the geometry of their waveguides can help reduce the number of parameters in the model. For instance, we used a physically inspired parametrization of the mode mixing matrices $M_{1,2}$ in the form of coupled waveguide lattices, which required only four real parameters per matrix, instead of nine, for the universal triangular MZI mesh parametrization. With this parametrization of $M_{1,2}$ and the same parametrization of the phase shift matrix, the total number of real parameters in our model is reduced from 30 to 20, which makes it significantly easier for the optimizer to find the optimal solution.
Technological advancements will also reduce the complexity of digital models. For example, the manufacturing of thermal phase shifters without mutual cross-interference eliminates all non-diagonal parameters $\alpha_{ij}$ from the crosstalk matrix $A$, resulting in N1N-1 real parameters, which are the diagonal elements of $A$, instead of $N-1\times N-1$ for $A$ with all nonzero elements. Alternatively, heaters can be designed to be straight and long relative to their width in order to simplify the model (7) of the phase shift layer with only two independent parameters (16), $\alpha_{0}$ and $\xi=d/r_{0}$, rather than $N-1\times N-1$ values of $\alpha_{ij}$ (see Appendix D).
Appendix H Machine learning based method for chip programming
@@
We also performed the learning of an artificial neural network on unitary matrices $U_{exp}$ measured on chip to reconstruct the mode mixing layers $M_{1,2}$ according to [40].
Both training and validation data sets unitaries, respectively, were taken from previously measured unitary matrices on the chip (see Fig. 4a). Learning was performed both on phase shifts - matrices dataset and currents - matrices dataset.
H.1 Phase shifts-matrices ($\vec{\varphi}$-$U$) dataset
The dataset contained $27$ pairs of unitary matrices and corresponding phase shifts $\vec{\varphi}$ obtained from the experimental calibration. This means that in such a dataset, we have knowledge of the phase shifts $\vec{\varphi}$ which were set on the chip and the corresponding measured unitary transformation $U_{data}(\vec{\varphi})$ of the chip. The neural network was trained on $27$ unitary matrices for which the fidelity between them and the calibration model is greater than $0.95$ (see Fig. 4a) and tested on the other $65$ measured unitary matrices.
The trained model was:
where $f_{rb}(\cdot)$ is a real border function defined in Appendix D, $A_{1,2}$ is two arbitrary complex matrices (not necessarily unitary), and $P(\vec{\varphi})=diag(e^{i\varphi_{1}},e^{i\varphi_{2}},e^{i\varphi_{3}},1)$ is a diagonal phase shift matrix. Matrices $A_{1,2}$ were considered to be general complex valued $4\times 4$ matrices and were parametrized by $2\times 4^{2}=32$ real parameters $\vec{p}$ each, which represent element wise parametrization of real and imaginary parts of $A_{1,2}$. The initial values of $\vec{p}$ were chosen randomly for each optimization run.
We used the Adaptive Moment Estimation optimization algorithm (Adam) on $5000$ epochs with mini batches containing $5$ random elements each. The Frobenius metric was chosen as the loss function, as in [40].
The results are presented in Fig. 10a. As can be seen from the results, the fidelity between the measured unitary matrices and predicted by the ML model is no less than $0.90$ with a peak around $0.99$, which indicates the accuracy of the revealed ML-based optical chip model. However, the phase shifts - matrices dataset require the knowledge of phase shifts information, which is obtained from the calibration procedure that provides the mode mixing elements $M_{1,2}$ itself. Therefore, performing another optimization procedure to determine the $A_{1,2}$ matrices, while already having the $M_{1,2}$ may seem somewhat redundant. In this regard, a more fundamental dataset of currents and corresponding unitary matrices for chips’ transformation is in high demand.
H.2 Currents-matrices ($\vec{x}$ - $U$) dataset
The dataset contained $60$ pairs of unitary matrices and the corresponding currents $\vec{x}$ set on the chip. Such scenario has more practical importance than the $\vec{\varphi}$ - $U$ dataset, as it contains solely primal raw data without any previous processing of it like calibration routine.
The neural network was trained on $60$ random unitaries from measured matrices on chip (see Fig. 4a) and tested on the other $36$ measured unitary matrices. In contrast to $\vec{\varphi}$ - $U$ dataset in $\vec{x}$ - $U$ dataset, we assume no calibration $\varphi(x)$ information and hence cannot choose $60$ unitaries with calibration model fidelity greater than $0.95$ as in the previous subsection.
The trained model was:
where $f_{rb}(\cdot)$ is a real border function, $A_{1,2}$ is two arbitrary complex matrices (not necessarily unitary), and $P(\vec{x})=diag(e^{i\varphi_{1}(\vec{x})},e^{i\varphi_{2}(\vec{x})},e^{i\varphi_{3}(\vec{x})},1)$ is a diagonal phase shift matrix with phase shifts from current dependence (7), resulting in $12$ real parameters: $3$ bias phase shifts $\vec{\varphi}_{0}$ and $9$ crosstalk matrix $A$ elements. Matrices $A_{1,2}$ were again considered to be general complex valued $4\times 4$ matrices and were parametrized by $32$ real parameters. The initial values for $\vec{p}$ were chosen at random for each optimization run, while an initial guess for the crosstalk matrix $\{\alpha_{ij}\}$ elements was taken according to the estimations discussed in the previous section (see Appendix D).
To train the network, the Adaptive Moment Estimation optimization algorithm was used on $6000$ epochs with mini batches containing $5$ random elements each. The Frobenius metric was chosen as the loss function as in [40].
The results are presented in Fig. 10b. As can be seen from the results, the fidelity between the measured unitary matrices and predicted by the trained ML model has a peak around $0.99$; however, some points are less than $0.90$, which proves the accuracy of the revealed ML-based optical chip model.
It is worth mentioning that in this experiment, training the artificial network on a currents-matrices ($\vec{x}$ - $U$) dataset to determine the ML-based optical chip model (34) aims to solve exactly the same task as the calibration of the phase shift layer. Both methods return the digital chip model as a function of the experimentally tunable parameters: electrical currents $\vec{x}$. This model enables the prediction of the unitary transformation of the chip at a given currents. However, despite sharing the same goal, these two methods differ in the experimental data they utilize. The calibration method requires all the output power from the current dependencies, whereas the ML method requires a set of unitary matrices measured on the chip. Under realistic experimental conditions, it may not always be easy to perform full optical chip tomography, in contrast to simple output power calibration measurements. At least, because if we use the same experimental resources (coherent laser radiation), full chip tomography would require injecting the radiation into two ports of the chip simultaneously and precisely varying their relative phase shift. However, only one input port was used in the calibration measurements. Both methods then use optimization over a set of real-valued parameters to minimize the corresponding loss function.

### v2

@@ Appendix E Unitary transformation of the optical chip measurement
where 2N1=72N-1=7 values of $\tilde{\gamma}_{ij}$ have zero values by construction. These are the angles of the $l_{1}$-th column and the $l_{1}$-th row of the reconstructed matrix $\tilde{U}$ on the chip. The fact that in the reconstructed unitary transformation $\tilde{U}$ of the target $N$-port interferometer, 2N12N-1 angles are set to be real is inherent to any tomography algorithm [47, 43, 44, 48]. This is due to the physical equivalence of $U$ and $\tilde{U}=P_{out}(\vec{\varphi_{out}})\times U\times P_{in}(\vec{\varphi_{in}})$, where $P_{in}(\vec{\varphi_{in}})$ and $P_{out}(\vec{\varphi_{out}})$ are diagonal matrices with pure phases $\vec{\varphi_{in}}$ and $\vec{\varphi_{out}}$, respectively. In other words, any tomography algorithm can reconstruct the unitary transformation of a linear optical interferometer only up to the input and output phases [47, 43, 44, 48].
As a final remark on the reconstruction procedure, the modules and the angles measurements are conducted completely independently, which does not provide any guaranty that the reconstructed matrix by (24) $\tilde{U}$ is unitary. Even in the case where one had performed the Sinkhorn-Knopp algorithm [46] to rescale the experimentally measured modules of $\tilde{U}$, the angles were still obtained using a separate interferometric procedure. However, if $\tilde{U}$ appeared to be nonunitary, it can be projected onto the closest matrix in the unitary subspace by, for example, polar decomposition or single value decomposition [49].
As, typically, optical power from all output ports of the chip is collected simultaneously each port by each photo detector, the number of separate measurements necessary to obtain the modules $|U_{ij}|$ for a $N\times N$-port interferometer is equal to $N$. However, angle measurement $\tilde{\gamma}_{ij}$ requires the collection of N1N-1 interference fringes, resulting in a total number of measurements for the unitary matrix $N\times N$ reconstruction of 2N12N-1.
In our experiment, we set $l_{1}=1$ for every measurement of $\tilde{U}$ on the chip, which makes the first row and the first column of any measured $\tilde{U}$ real. However, because the measured unitary matrix of the target interferometer $\tilde{U}$ is in the form of a real first row and first column, the simulated matrix $U_{sim}$, generated by the chips’ digital model, should be converted to this same form before their comparison, since the matrix fidelity (9) measure is sensitive to such deviations, while these unitary matrices are physically equivalent. This type of unitary matrix $U_{sim}=\rho_{ij}\exp(i\varphi_{ij})$ transformation can be simply performed by multiplying it on both sides by diagonal unitary matrices:
@@
where $\Gamma=diag(e^{i\gamma_{j}})$ and $\Omega=diag(e^{i\omega_{i}})$, the elements of which are equal to $\gamma_{j}=-\varphi_{1j}+\varphi_{11}$ and $\omega_{i}=-\varphi_{i1}$. After multiplying by $\Gamma$ and $\Omega$, the unitary matrix $U_{sim}$ transforms to $\tilde{U}_{sim}=\rho_{ij}\exp(i\tilde{\varphi}_{ij})$ with zero angles in the first row and first column by construction: $\tilde{\varphi}_{1j}\equiv\tilde{\varphi}_{i1}\equiv 0$.
More generally, any $N\times N$ unitary matrix $U_{ij}=\rho_{ij}\exp(i\varphi_{ij})$ can always be multiplied on both sides by diagonal $N\times N$ unitary matrices $\Gamma=\text{diag}(e^{i\gamma_{j}})$ and $\Omega=\text{diag}(e^{i\omega_{i}})$ with appropriate angles $\gamma_{j}$ and $\omega_{i}$ such that the resulting unitary matrix $\tilde{U}_{ij}=\rho_{ij}\exp(i\tilde{\varphi}_{ij})$ has real $I$-th row and $J$-th column: $\tilde{\varphi}_{iJ}\equiv\tilde{\varphi}_{Ij}\equiv 0$. The necessary angles $\gamma_{j}$ and $\omega_{i}$ are obtained from:
@@
Appendix F Phase shifts-matrices ($\vec{\varphi}$-$U$) dataset
The data set contained $27$ pairs of unitary matrices and the corresponding phase shifts $\vec{\varphi}$ obtained from the experimental calibration. This means that, in this data set, we have the knowledge of the exact phase shifts $\vec{\varphi}$ that were set on the chip and the corresponding measured unitary transformation $U_{data}(\vec{\varphi})$ of the chip. The neural network was trained on $27$ unitary matrices for which the fidelity between them and the calibration model is greater than $0.95$ (see Fig. 4a) and tested on the other $65$ measured unitary matrices.
The trained model was:
where $f_{rb}(\cdot)$ is a real border function defined in Appendix D, $A_{1,2}$ is two arbitrary complex matrices (not necessarily unitary), and $P(\vec{\varphi})=diag(e^{i\varphi_{1}},e^{i\varphi_{2}},e^{i\varphi_{3}},1)$ is a diagonal phase shift matrix. The matrices $A_{1,2}$ were considered as general complex valued $4\times 4$ matrices and were parameterized by $2\times 4^{2}=32$ real parameters $\vec{p}$ each, representing the parameterization of the real and imaginary parts of $A_{1,2}$. The initial values of $\vec{p}$ were chosen at random for each optimization run.
We used the Adaptive Moment Estimation optimization algorithm (Adam) with $5000$ epochs with mini batches containing $5$ random elements each. The Frobenius metric was chosen as the loss function, as in [40].
The results are presented in Fig. 5a. As can be seen from the results, the fidelity between the measured unitary matrices and predicted by the ML model is no less than $0.90$ with a peak around $0.99$, which indicates the accuracy of the revealed ML-based optical chip model. However, the phase shifts - matrices data set requires exact knowledge of the phase shifts information, which is obtained from the calibration procedure that provides the mode mixing elements $M_{1,2}$ itself. Therefore, performing another optimization procedure to determine the $A_{1,2}$ matrices, while already having the $M_{1,2}$ may seem somewhat redundant. In this regard, a more fundamental dataset of currents and corresponding unitary matrices for chips’ transformation is in high demand.
Appendix G Currents-matrices ($\vec{x}$ - $U$) dataset
The data set contained $60$ pairs of unitary matrices and the corresponding currents $\vec{x}$ set on the chip. This scenario is of more practical importance than the $\vec{\varphi}$ - $U$ data set, since it contains only raw primal data without any previous processing, such as the calibration routine.
The neural network was trained on $60$ random unitary matrices from measured matrices on the optical chip (see Fig. 4a) and tested on the other $36$ measured unitary matrices. In contrast with data set $\vec{\varphi}$ - $U$ in data set $\vec{x}$ - $U$, we assume that there is no calibration information $\varphi(x)$ and therefore cannot choose $60$ unitary matrices with calibration model fidelity greater than $0.95$ as in the previous subsection.
The trained model was:
where $f_{rb}(\cdot)$ is a real border function, $A_{1,2}$ is two arbitrary complex matrices (not necessarily unitary), and $P(\vec{x})=diag(e^{i\varphi_{1}(\vec{x})},e^{i\varphi_{2}(\vec{x})},e^{i\varphi_{3}(\vec{x})},1)$ is a diagonal phase shift matrix with phase shifts from current dependence (7), resulting in $12$ real parameters: $3$ bias phase shifts $\vec{\varphi}_{0}$ and $9$ crosstalk matrix elements $A_{ij}$. The matrices $A_{1,2}$ were again considered to be general complex-valued matrices $4\times 4$ and were parameterized by $32$ real parameters. The initial values for $\vec{p}$ were chosen at random for each optimization run, while an initial guess for the crosstalk matrix elements $\{\alpha_{ij}\}$ was taken according to the estimations discussed in the previous section (see Appendix D).
To train the network, the Adaptive Moment Estimation optimization algorithm was used with $6000$ epochs with mini batches containing $5$ random elements each. The Frobenius metric was chosen as the loss function as in [40].
Appendix H Broadband optical switching details
@@
Optical switches were realized exploiting the chips’ models at several wavelengths ranging from $915$ to $975$ nm. At each wavelength, the digital chip model was reconstructed using the calibration method. The reconstructed mode mixing matrices $M_{1,2}$, bias phase shift vectors $\vec{\Phi}_{0}$ and crosstalk matrices $A$ for each wavelength are shown in Fig. 11.
@@
As shown in Fig. 11 the mode mixing matrices $M^{\lambda}_{1,2}$ are slightly different for each wavelength $\lambda$ preserving the property $M^{\lambda}_{1}\approx M^{\lambda}_{2}$. The bias phase shift vectors $\vec{\Phi}_{0}$ vary with wavelength $\lambda$ more significantly, but always in accordance with the actual geometry of the target interferometer: $\varphi_{1}^{\lambda}\approx 0$ and $\varphi_{2}^{\lambda}\approx\varphi_{3}^{\lambda}$ as discussed in the Appendix D. Meanwhile, the crosstalk matrices $A^{\lambda}$ for each wavelength $\lambda$ are approximately the same. This is because the crosstalk matrix $A$ describes the physical heaters (and is completely determined by them) and not the coherent radiation.
@@
The necessary three phase shifts for implementing a particular port-to-port optical switch were obtained by an optimization procedure run on a standard PC (Python scipy.minimize method). Once the required phase shifts were determined from the digital model of the optical chip, they were applied to the actual device. Coherent radiation of a given wavelength was injected into the specific input port of the device. The optical power distribution at the output ports was then measured using photodiodes and compared with the target values. The comparison between the switching matrices $U$ and $V$ was made with respect to the variation in the amplitude fidelity measure:
@@
An example of simulating coherent light with a wavelength of $975$ nanometers propagating through a target interferometer, illustrating all possible port-to-port optical switch configurations, is shown in Fig. 12.
Appendix I Scaling of the proposed calibration procedure
The scalability of the proposed approach for programming a reconfigurable integrated interferometer is a crucial consideration that needs to be carefully addressed. First, in our experiment, the tunable integrated interferometer consists of two passive mode mixing blocks $M_{1,2}$ and an active phase shift layer $P_{1}(\vec{\varphi}_{1})$ between them. The phase shift layer consists of three thermos-optical phase shifters. Therefore, our optical chip has only three tunable parameters and cannot be considered a $4\times 4$ universal photonic device. However, if two additional phase shift layers $P_{2}(\vec{\varphi}_{2})$ and $P_{3}(\vec{\varphi}_{3})$ were cascaded between the two additional constant mode mixing blocks $M_{3}$ and $M_{4}$ and added to our device, the resulting interferometer would become the universal interferometer [9, 31, 10].
@@ Appendix I Scaling of the proposed calibration procedure
This calibration will result in the exact phase-current dependence for the first phase shift layer, as well as in the explicit form of the first mode mixing block $M_{1}$ and the effective mode mixing block $M_{2}$:
@@
Figure 10: a) A model of an interferometer with two phase shift layers ($P_{1}$ and $P_{2}$) and three mode mixing layers ($M_{1}$, $M_{2}$ and $M_{3}$) can be consequently treated as b) interferometer with one phase shift layer $P_{2}$ cascaded between two mode mixing layers $\tilde{M}_{1}$ and $M_{3}$, and as c) interferometer with one phase shift layer $P_{1}$ cascaded between two mode mixing layers $M_{1}$ and $\tilde{M}_{2}$. Mode mixers, $M_{1}$ and $M_{3}$, and phase shift layers, $P_{1}$ and $P_{2}$, can be evaluated from two independent calibration routines described in the main text. The mode mixing block $M_{2}$ can then be implicitly determined from the auxiliary $\tilde{M}_{1}$ and $\tilde{M}_{2}$.
This calibration will result in the exact phase-current dependence for the second phase shift layer, as well as in the explicit form of the third mode mixing block $M_{3}$ and the effective mode mixing block $M_{1}$:
@@
which, again, just as $\tilde{M}_{2}$ implicitly contains information about the actual $M_{2}$. However, we know the explicit forms of $M_{1}$ and $P_{1}(\vec{\varphi}_{1}^{(0)})$ (which is the first phase shift layer with zero currents). This enables us to determine the actual second mode mixing block $M_{2}$ as follows:
@@
The expressions (34) and (35) for the calculation of $M_{2}$ are independent, since the entities $M_{1}$, $P_{1}^{(0)}$, $\tilde{M}_{2}$ and $M_{3}$, $P_{2}^{(0)}$, $\tilde{M}_{1}$ were obtained from two independent experimental calibration procedures for the first (32) and second (33) phase shift layers $P_{1}(\vec{\varphi}_{1}(\vec{X}_{1}))$ and $P_{2}(\vec{\varphi}_{2}(\vec{X}_{2}))$, respectively. The two independent evaluations of $M_{2}$ must be equivalent and can be used to validate the entire optical chip digital model. Figure 10 schematically illustrates the procedure described above for programming an interferometer with two phase shift layers.
@@
The necessary number of measurements for the calibration of a phase shift layer in a general $N\times N$ interferometer is $N\times(N-1)$ as a product of the input port number and the number of heaters. This scaling factor is quadratic with respect to the number of modes. The number of real parameters required for optimization is also quadratic in the number of modes.
In other words, there are two ways to scale an interferometer: by increasing its length, or by increasing the width or the number of modes it contains. The first method is simpler and involves a linear increase in the number of measurements and amount of processing required. However, scaling by increasing the number of modes grows quadratically with both the number of measurements and the number of parameters of the model that need to be optimized.
While the number of measurements cannot be decreased by increasing the number of modes, $N$, the number of parameters in the model can be reduced by carefully adjusting the model. Therefore, finding possible methods to reduce the number of model parameters is of great importance. For example, using prior knowledge of the structure of actual interferometers and the geometry of their waveguides can help reduce the number of parameters in the model. For example, we used a physically inspired parametrization of the mode mixing matrices $M_{1,2}$ in the form of coupled waveguide lattices, which required only four real parameters per matrix, instead of nine, for the universal triangular MZI mesh parametrization. With this parameterization of $M_{1,2}$ and the same parameterization of the phase shift matrix, the total number of real parameters in our model is reduced from 30 to 20, making it significantly easier for the optimizer to find the optimal solution.
Technological advancements will also reduce the complexity of digital models. For example, manufacturing of thermal phase shifters without mutual cross-interference eliminates all non-diagonal parameters $\alpha_{ij}$ from the crosstalk matrix $A$, resulting in N1N-1 real parameters, which are the diagonal elements of $A$, instead of $N-1\times N-1$ for $A$ with all nonzero elements. Alternatively, heaters can be designed to be straight and long relative to their width to simplify the model (7) of the phase shift layer to one with only two independent parameters (17), $\alpha_{0}$ and $\xi=d/r_{0}$, instead of $N-1\times N-1$ values of $\alpha_{ij}$ (see Appendix D).